OpenStudy (anonymous):

Determine the quadratic equation for the parabola: minimum value: -5; x-intercepts: -2 and 3

OpenStudy (anonymous):

Please help! So stuck!

OpenStudy (amoodarya):

y=a(x-3)(x-(-2)) so put axis of symmetry into it and find a y=-5 when put x of axis of symmetry

OpenStudy (anonymous):

I don't understand :(

OpenStudy (amoodarya):

y=a(x-x1)(x-x2) axis of symmetry is x1 +x2 x= -------- 2

OpenStudy (anonymous):

I have to get the answers like this, f(x) = 4/5x2 + 2/5x - 48/5 f(x) = 2x2 + 4x - 48 f(x) = 2/5x2 + 4/5x - 48/5 f(x) = 5x2 + 4x - 48

OpenStudy (anonymous):

These are the answer im given and I am not getting how to get these

OpenStudy (amoodarya):

x1=-2 x2=+3 so axis of symmetry is -2+3 1 x=------=---- 2 2 so put x=1/2 to y=a(x+2)(x-3) this is minimum value -5=a(1/2+2)(1/2-3) -5=a(5/2)(-5/2) so a=4/5

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

Do I do 4/5(5/2(-5/2) and distribute?

OpenStudy (amoodarya):

yes solve for a

OpenStudy (anonymous):

Doesn't make sense I got 2 and -2

OpenStudy (amoodarya):

\[y=a(x+2)(x-3) \rightarrow put\x=\frac{ 1 }{2 } \\-5=a(\frac{ 1 }{2 }+2)(\frac{ 1 }{2 }-3) \\-5=a(\frac{ 5 }{2 })(\frac{ 1-6 }{2 })\\-5=a(\frac{ 5 }{2 })(\frac{ -5 }{2 })\\-5=a(\frac{- 25 }{4 })\\ \frac{ -5 }{ 1 }=a(\frac{- 25 }{4 })\\ \frac{ -5*4 }{ -25 }=a\\a=\frac{ 4 }{ 5 }\]